A meat market manager mixes hamburger having 30% fat content with ground round that has 10% fat content in order to obtain 400 pounds of meat with 25% fat content. How much beef of each type should the manager use?
Thank you

Answer by checkley77(12844)

.30x+.10(400-x)=.25*400
.30x+40-.10x=100
.20x=100-40
.20x=60
x=60/.20
x=300 pounds of 25% hamburger is used.
400-300=100 pounds of hamburgers is used.
Proof:
.30*300+.10*100=.25*400
90+10=100
100=100